OpenStudy (anonymous):

..Theresa is eight years younger than Sian. If the sum of their ages is 34, how old is Therese?

OpenStudy (anonymous):

13....?

OpenStudy (anonymous):

13 plus 8 = 21 not 34

OpenStudy (anonymous):

yaa bt 13 +21 is 34......

OpenStudy (anonymous):

Let x = Sian's age x - 8 = Theresa's age x + (x - 8) = 34 Solve for x.

OpenStudy (anonymous):

let sian's age be 'x' and teresa's age be 'x-8' x-8+x = 34 2x-8 = 34 2x = 34+8 x = 42/2 x = 21 so,sian's age=21 teresa's age = 21-8 = 13

OpenStudy (anonymous):

thanks mate...perfect...

OpenStudy (anonymous):

always welcome

OpenStudy (anonymous):

no issue..:)